# Design principles of the cytotoxic CD8+ T-cell response

### v2

@@ Results
To align responses with the observed macro-dynamic trends, subsequent analyses consider near-elbow designs $\vec{\theta}\in\Theta(\lambda^{*})$ with their infection-signal sensitivity reduced to one-tenth of their original value: for $\psi_{\text{inf}}\mapsto 0.1\,\psi_{\text{inf}}$. This ensemble of tuned designs is indicated by $\tilde{\Theta}(\lambda^{*})$. See SI Appendix, Figs. S11 and S12 for extensive model parameter sensitivity analyses with this ensemble.
@@
CD8+ T-cells are central to cancer surveillance and elimination of mutated pre-cancerous cells. We therefore asked how the tuned elbow-optimal designs $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$ identified for infectious threats (Fig. 4B, Fig. 5) respond to cancer, which we modeled as a growing, low-immunogenic tumor that kills healthy cells as it expands (Fig. 6A; SI Appendix).
Similar to the case of infection control, a clearance-toxicity trade-off arises, but it is sharpened in cancer because many tumor antigens are only marginally distinct from self. Even high-affinity recognition risks cross-reactivity and bystander damage. Tuned near-elbow designs incur high total harm for low-immunogenic tumors (Fig. 6B,C). As tumor immunogenicity increases, harm declines due to decreases in cross-reactive toxicity and increases in clearance (Fig. 6C).
@@
TCR engineering raises an additional nuance: increasing TCR-tumor affinity often increases affinity to self [69], so that increasing tumor immunogenicity is difficult. To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged (SI Appendix, Fig. S10). We quantify affinity or reactivity of a cell by the ratio of the maximum healthy cell counts $S_{\text{max}}$ to the number of target cells required for half-maximal T-cell response (EC50), i.e., $K_{I}^{-1}S_{\text{max}}$ for tumor cells, and $K_{S}^{-1}S_{\text{max}}$ for self targets. Provided that tumor immunogenicity ($\kappa=K_{I}^{-1}K_{S}$) is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig. S10A) mirrors its dependence on immunogenicity in Fig. 6B. Nevertheless, these cross-reactive regimes (Fig. S10B) yield higher total harm than scenarios where self-interaction fixed so that increasing tumor antigenicity increases immunogenicity (Fig. 6C). This highlights the cost of affinity-driven cross-reactivity in TCR engineering for cancer therapy [70, 71, 72].
@@ Discussion
In this work, we modeled the CD8+ T-cell response as a feedback-controlled program that integrates three cues—instantaneous antigen, cumulative infection-derived harm, and cumulative response-derived harm—to regulate activation, expansion, differentiation, and contraction of T-cells. Modulation of response proportional to the antigen signal and the time-integrated harm signals mirrors proportional-integral-derivative (PID) control of engineered systems, a strategy known for its stable and robust steering of dynamical processes toward desired outcomes [80].
@@
Exploring an ensemble of biologically motivated controllers (designs) and optimizing for pathogen clearance while minimizing immunopathology exposed a Pareto trade-off: a reduction in infection harm generally leads to an increased collateral damage to healthy tissues. This pattern persisted across diverse pathogenic challenges defined by their immunogenicity and within-host basic reproduction number.
@@
Three design principles emerged: (i) Pre-programmed (baseline-tuned) designs help restrain autoimmunity and slowly proliferating threats but are inefficient against fast-proliferating infections or metastatic tumors; effective control there requires signal integration and feedback. (ii) Among signal sources, antigen and response-harm dominate; sensitivity to infection-harm contributes comparatively little once CD8+ effector T-cell expansion is underway. (iii) Feedback from response-harm acts as an anti-inflammatory brake, enabling strong antigen-driven responses while limiting collateral damage when the response overshoots—a characteristic consistent with endogenous effector programs.
@@
We posited that effector programs should attempt to solve a multi-objective problem to counter diverse pathogenic challenges. Interestingly, we found that a Pareto-optimal design that suppresses responses to frequent, self-like injuries while clearing recognizable, replicating (e.g., acute) infections mirrors the endogenous T-cell programs in vertebrates: strongly antigen-driven with an anti-inflammatory brake, and conservative cell-fate transition speeds that prevent runaway effector responses.
We assumed that, although memory formation is essential, its immediate impact on primary effector magnitude is modest: $5-10\%$ of peak effector numbers [81, 16]. Memory “quality” (longevity, proliferative capacity) depends on differentiation timing and the avidity of the selected clones [82, 15, 21, 83]. To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall. An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the host’s evolutionary survival pressures that could have shaped the effector program.
@@
Using a minimal tumor model, we showed how the design objective informs rational immunotherapy protocols. First, it suggests that infused T-cell dosage should depend on tumor immunogenicity—balancing clearance against collateral damage, as previously emphasized in the literature [66, 71, 68]. Second, treating the design parameters as engineering targets yields testable strategies to improve efficacy and safety, which is highly desirable [84, 85, 86]. For moderately immunogenic tumors, promoting effector output by raising baseline rates of activation/proliferation or increasing antigen and infection sensitivities, while tempering the anti-inflammatory brake, can produce robust responses.
@@
Key questions follow from our study. Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs [88, 89]? Endogenous immune responses are polyclonal [90]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses? Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them? Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
II Data and code availability
@@
Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune.

### v3

@@ III Results
To align responses with the observed macro-dynamic trends, subsequent analyses consider near-elbow designs $\vec{\theta}\in\Theta(\lambda^{*})$ with their infection-signal sensitivity reduced to one-tenth of their original value: for $\psi_{\text{inf}}\mapsto 0.1\,\psi_{\text{inf}}$. This ensemble of tuned designs is indicated by $\tilde{\Theta}(\lambda^{*})$.
@@
CD8+ T-cells are central to cancer surveillance and elimination of mutated pre-cancerous cells. We therefore asked how the tuned elbow-optimal designs $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$ identified for infectious threats (Fig. 4B, Fig. 5) respond to cancer, which we modeled as a growing, low-immunogenic tumor that kills healthy cells as it expands (Fig. 6A; SI Appendix). We quantified harm as the total number of healthy cells lost, combining cell death caused directly by tumor expansion (replacing healthy cells) and collateral killing mediated by the immune response (Fig. 6B,C).
Similar to the case of infection control, a clearance-toxicity trade-off arises, but it is sharpened in cancer because many tumor antigens are only marginally distinct from self. Even high-affinity recognition risks cross-reactivity and bystander damage. Tuned near-elbow designs incur high total harm for low-immunogenic tumors (Fig. 6B, C). As tumor immunogenicity increases, harm decreases due to reduced cross-reactive toxicity and enhanced tumor clearance (Fig. 6C).
@@
TCR engineering raises an additional nuance: increasing TCR-tumor affinity often increases affinity to self [70], so that increasing tumor immunogenicity is difficult. To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged ($\kappa=K_{I}^{-1}K_{S}=\text{constant}$; SI Appendix, Fig. S9). We quantify affinity or reactivity of a cell by the ratio of the maximum healthy cell counts $S_{\text{max}}$ to the number of target cells required for half-maximal T-cell response (EC50), i.e., $K_{I}^{-1}S_{\text{max}}$ for tumor cells, and $K_{S}^{-1}S_{\text{max}}$ for self targets. Provided that tumor immunogenicity $\kappa$ is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig. S9A) mirrors its dependence on immunogenicity in Fig. 6B. Nevertheless, these cross-reactive regimes (Fig. S9B) yield higher total harm than scenarios where self-interaction is fixed so that increasing tumor antigenicity increases immunogenicity (Fig. 6C). This highlights the cost of affinity-driven cross-reactivity in TCR engineering for cancer therapy [71, 72, 73].
@@ IV Discussion
In this work, we modeled the CD8+ T-cell response as a feedback-controlled program that integrates three cues (i.e., instantaneous antigen, time-integrated infection-derived harm, and time-integrated response-derived harm) to minimize cumulative cell death during an immune challenge. Modulation of response proportional to the antigen signal and the time-integrated harm signals mirrors proportional-integral-derivative (PID) control of engineered systems, a strategy known for its stable and robust steering of dynamical processes toward desired outcomes [81].
@@
Exploring an ensemble of biologically motivated controllers (designs) and optimizing for pathogen clearance while minimizing immunopathology exposed a Pareto trade-off: a reduction in infection harm generally leads to an increased collateral damage to healthy tissues. This pattern persisted across diverse immune challenges defined by their immunogenicity and within-host basic reproduction number.
@@
Three design principles emerged from our analyses: (i) Pre-programmed (baseline-tuned) designs help restrain autoimmunity and slowly proliferating threats but are inefficient against fast-proliferating infections or metastatic tumors; effective control there requires signal integration and feedback. (ii) Among signal sources, antigen and response-harm dominate; sensitivity to infection-harm contributes comparatively little once CD8+ effector T-cell expansion is underway. (iii) Feedback from response-harm acts as an anti-inflammatory brake, enabling strong antigen-driven responses while limiting collateral damage when the response overshoots—a characteristic consistent with endogenous effector programs [29, 28]. Table 1 further summarizes the characteristics of our inferred optimal designs and relates them to supporting evidence from endogenous effector programs.
@@
We posited that effector programs should attempt to solve a multi-objective problem to minimize harm (cell death) inflicted by diverse pathogenic challenges. Interestingly, we found that a Pareto-optimal design that suppresses responses to frequent, self-like injuries while clearing recognizable, replicating (e.g., acute) infections mirrors the endogenous T-cell programs in vertebrates: strongly antigen-driven with an anti-inflammatory brake, and conservative cell-fate transition speeds that prevent runaway effector responses. We note that, beyond minimizing harm due to cell death, additional factors may also shape the objectives optimized by effector T-cell programs, including the metabolic costs of effector activity and cytokine production, the cumulative burden of inflammation, and the long-term value of generating memory cells. Although the relative importance of these contributions remains unclear, they could be incorporated and analyzed within the same Pareto-optimization framework we introduced in this work.
Memory formation is a key feature of CD8+ T-cell immunity, but here we focus on acute infection control and therefore do not model memory expansion or recall explicitly. In endogenous responses, immunological memory arises in part from the differentiation of a small fraction of effector cells ($\sim 5-10\%$ of the peak response [82, 16]) into effector-memory cells, effectively removing them from the active response. In our model, this loss of effector cells is captured by the contraction step, which can represent either effector apoptosis or differentiation into memory. By aggregating these processes into a single contraction term, we reduce degeneracy in the design parameters while retaining interpretability. Nonetheless, memory “quality” depends on differentiation timing and clone avidity [83, 15, 21, 84], which would be relevant for the strength and durability of protection upon secondary immune challenges. Extending our framework to an explicit allocation trade-off, where diverting cells to memory reduces primary effector output but seeds future protection, will be an important next step toward linking control designs to memory formation and the evolutionary pressures shaping effector programs.
@@
Using a minimal tumor model, we showed how the design objective informs rational immunotherapy protocols. First, it suggests that infused T-cell dosage should depend on tumor immunogenicity—balancing clearance against collateral damage, as previously emphasized in the literature [67, 72, 69]. Second, treating the design parameters as engineering targets yields testable strategies to improve efficacy and safety, which is highly desirable [85, 86, 87]. For moderately immunogenic tumors, promoting effector output by raising baseline rates of activation/proliferation or increasing antigen and infection sensitivities, while tempering the anti-inflammatory brake, may produce robust responses.
@@
Key questions follow from our study. Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs [89, 90]? Endogenous immune responses are polyclonal [91]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses? Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them? Would explicitly modeling interactions between CD8+ T-cells and other immune compartments, such as regulatory T-cells and innate immune system, enhance the robustness of cytotoxic responses across different pathogenic challenges? Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
V Materials and Methods
@@
Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune. Previously published data from refs. [9, 63] were used for this work.
Intra-host infection dynamics. We model a population of healthy, susceptible cells $S(t)$, initially consisting of $S_{\text{max}}$ identical cells. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs. 1A,B, S1). The dynamics of the susceptible and infected cells follow,
where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [92, 93, 94]. The functional form of the last terms in these equations is motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix). Note that these equations describe dynamics during the immune challenge and exclude the slow birth-death processes that homeostatically maintain the susceptible cells in the absence of infection.
Stochastic T-cell response model. We model the CD8+ T-cell response to an immune challenge as a time-inhomogeneous birth-death process coupled to deterministic dynamics of susceptible and infected cells. Beginning with an initial population of monoclonal naïve T-cells, we track individual T-cell lineages through recruitment of naïve cells to interact with antigen-presenting cells (APCs; $N\to\textbf{APC}\cdot N$), activation ($\textbf{APC}\cdot N\to N^{*}$), differentiation to effector ($N^{*}\to E$), expansion ($E\to 2E$), and contraction ($E\to\emptyset$). Naïve T-cells are recruited to APCs at a rate modulated by infection harm signals, and upon sustained engagement, become activated and undergo 2-3 burst-like divisions largely independent of continued antigenic input. During this early proliferative window, each activated cell commits to an effector or memory fate according to the first arrival of a time-inhomogeneous Poisson process with a signal-dependent rate; the number differentiating into effectors is drawn as a binomial random variable, with remaining cells adopting a memory phenotype. Effector cells subsequently divide and die at rates governed by time-varying antigen and harm signals. Proliferation is capped by a maximum division number [11, 7, 9, 10], which we set to 15 rounds, beyond which cells undergo terminal arrest (see SI Appendix for details).
In our model, inter-lineage variability arises solely from stochastic differences in transition timing, capturing with prior observations [17, 18, 16, 19, 20, 15, 11, 21]. To quantify this stochasticity and demonstrate robustness, results are averaged over five independent full-simulation replicates, with variability reported. The coefficient of variation averaged over immune challenges is small (C.V. = 0.05; Fig. S12), indicating that macroscopic quantities such as total harm and effector peak are robust to stochasticity in T-cell response dynamics. We also characterize sensitivity to simulation parameters (Figs. S10, S11), which vary across hosts and infection scenarios.
Infection and response signals. T-cell activation relies on the integration of antigen and harm signals. Cell deaths can induce harm in the host and produce pro- or anti-inflammatory signals, depending on the underlying cause. We simplify this by distinguishing infection-induced and response-induced harms, which in a cumulative form, generate signals $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. The instantaneous level of presented antigens, proxied by the infected-cell burden $I(t)$, defines a separate antigen signal $\sigma_{\text{Ag}}(t)$. Together, scaled by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
where $h_{\text{inf}}(s)$ is the number of cells killed by the pathogen (infection), and $h_{\text{res}}(s)$ denotes the number of cells killed by the effector T-cell response at time $s$ (eq. 3). The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).